what is wave number?

Grade:12th Pass

12 Answers

Kiran Allada
8 years ago
wave number (k) is used to describe a sinosoidal wave.
y(x,t)=Asin(kx-wt)  where k=2pi/lambda
karthik vinjarapu
8 years ago
 
wave number (k) is used to describe a sinosoidal wave.
y(x,t)=Asin(kx-wt)
Gavvala Ganesh
8 years ago
In Clark's processslaked lime, Ca(OH)2 is added to temporary hard water then the water will become as soft water it is one of the process to hardness of water.
SHANMUKESHWAR
8 years ago
=Asin(kx-wt) where k=2pi/lambda(x,t)wave number (k) is used to describe a sinosoidal wave. y
Gavvala Ganesh
8 years ago
 
wave number (k) is used to describe a sinosoidal wave.
y(x,t)=Asin(kx-wt)  where k=2pi/lambda
 
 
 
KUNCHAM SAMPATH
8 years ago
 
wave number (k) is used to describe a sinosoidal wave.
y(x,t)=Asin(kx-wt)  where k=2pi/lambda
pa1
8 years ago
=Asin(kx-wt) where k=2pi/lambda (x,t) wave number (k) is used to describe a sinosoidal wave. y
pa1
8 years ago
 
wave number (k) is used to describe a sinosoidal wave.
y(x,t)=Asin(kx-wt)  where k=2pi/lambda
venky
8 years ago
the number of waves in a unit distance.
SAI SARDAR
8 years ago
Dear Prapulpodisetti,
         Wave number is the the number of wavelengths per unit distance is called wave number.
                        k=2pi/lamda
Gowri sankar
8 years ago
MR PRAPUL,
                 Dear Prapulpodisetti,
         Wave number is the the number of wavelengths per unit distance is called wave number.
                        k=2pi/lamda
Prabhakar ch
8 years ago
the number of waves in a unit distance